OpenStudy (anonymous):

Can someone please explain how to make x the subject of the equation please? x²/5 = w +y

OpenStudy (anonymous):

v/4x = w/y v=4xw/y vy=4xw 4xw=vy 4x=vy/w x=vy/4w

OpenStudy (anonymous):

where di you get the v from?

OpenStudy (anonymous):

no this is an example (v/4x = w/y) thats how you solve it and make x the subject.

OpenStudy (anonymous):

oh ok thanks (: